# PsyDI: Towards a Personalized and Progressively In-depth Chatbot for Psychological Measurements

### v3

@@ Abstract
In the field of psychology, traditional assessment methods, such as standardized scales, are frequently critiqued for their static nature, lack of personalization, and reduced participant engagement, while counseling evaluations are often inaccessible to the general public. The complexity of quantifying psychological traits further limits these methods. Despite advances with large language models (LLMs), many still depend on single-round Question-and-Answer interactions. To bridge this gap, we introduce PsyDI, a personalized and progressively in-depth chatbot designed for psychological measurements, exemplified by its application in the Myers-Briggs Type Indicator (MBTI) framework. PsyDI leverages user-related multi-modal information and engages in customized, multi-turn interactions to provide personalized, easily accessible measurements, while ensuring precise MBTI type determination. To address the challenge of unquantifiable psychological traits, we introduce a novel training paradigm that involves learning the ranking of proxy variables associated with these traits, culminating in a robust score model for MBTI measurements. The score model enables PsyDI to conduct comprehensive and precise measurements through multi-turn interactions within a unified estimation context. Through various experiments, we validate the efficacy of both the score model and the PsyDI pipeline, demonstrating its potential to serve as a general framework for psychological measurements. Furthermore, the online deployment of PsyDI has garnered substantial user engagement, with over 3,000 visits, resulting in the collection of numerous multi-turn dialogues annotated with MBTI types, which facilitates further research. The source code for the training and web service components is publicly available as a part of OpenDILab at: https://github.com/opendilab/PsyDI
@@ 2 Problem Formulation
We initiate our analysis by modelling the measurement of psychological indicators based on the characteristics of MBTI, as detailed in the Appendix A. We conceptualize the interaction between LLM agents and the test taker (user) as a Markov Decision Process (MDP) (Sutton & Barto, 2018). This formulation allows us to systematically optimize the sequence of inquiries posed to the user, with the aim of enhancing the accuracy, efficiency and engagement of MBTI evaluations. Formally, the MDP is articulated as =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\}$. The states $s$ in the observable state space $\mathcal{S}$ are defined as $s=\{p_{1},\ldots,p_{n}\}$, where $p_{i}$ is the $i$-th textual statement of the user (e.g. I like to talk to different people and share my stories). Additionally, the state space includes a special terminated state $s_{\text{exit}}$ representing the user’s decision to exit the evaluation process.
@@ 3 PsyDI Progressively In-depth Pipeline
In this section, we present the motivation and detailed design of the PsyDI pipeline. The task of collecting quantifiable MBTI labels for machine learning applications presents a significant challenge. Nevertheless, we can draw upon the methodologies employed in classical psychological scales, which often include specific quantitative metrics to assess the test-taker’s during testing. In line with this approach, we have devised an MBTI profile that dynamically summarizes and records the MBTI predilection derived from the test-taker’s statements, as described in Section 3.1. This profile is continuously updated with new self-reported statements or feedback obtained from the summary of the multi-turn chat, thereby refining the MBTI profile in real-time. Leveraging this design, the remainder of PsyDI is structured into two distinct phases: statement selection and multi-turn chat. The former is dedicated to identifying statements with the greatest potential for questioning (Section 3.2). Then, the multi-turn chat phase is designed to engage users in iterative questions based on the selected statement, fostering an environment that encourages them to provide statements that accurately reflect their cognitive processes while also reducing the chance that they might stop participating (Section 3.3). Finally, PsyDI will ascertain the user’s MBTI upon achieving a confident estimation. The entire workflow of PsyDI is illustrated in Figure 2. For more details on the pipeline, please refer to Appendix D.
@@ 3.3 Multi-turn Chat
Although LLMs perform exceptionally well in single-round, goal-oriented dialogues, it encounters difficulties in maintaining a coherent focus around a single objective when faced with a dozen rounds of multiple questions. To address this limitation, we have devised a hybrid approach that leverages LLMs’ strengths by structuring interactions into modules of three to five rounds, each centered on a specific goal. Within these modules, their capabilities are harnessed to progressively delve into questions related to MBTI, ensuring a step-by-step deepening of the exploration. The summarization of each module’s dialogue then contributes new insights to the pipeline, facilitating a cumulative and systematic advancement of the overall structure.
@@
For the first challenge, to lead users to reveal their underlying motivations, we utilize multiple-choice questions where each option corresponds to a specific cognitive preference. By presenting such options, users can select the one that best aligns with their own, thereby avoiding ambiguous expressions and implicitly guiding them to reflect on their cognitive tendencies. For questions requiring confirmation, PsyDI employs forced-choice questions, compelling users to identify their most likely cognitive preferences. Also, PsyDI accepts free-form answers, allowing users to provide responses beyond the preset options. The prompt for this phase is list in Appendix C. For the second challenge, which is to keep user’s interesting. We have designed a multimodal input system that allows users to input their preferred music and images, which are then converted into text to form a user statement. Additionally, our questioning framework is designed with a three-phase thought chain. First, we analyze the user statement, then we infer the user’s underlying thought patterns, and finally, we derive the next question based on these patterns. This approach ensures that the entire questioning process revolves around the user’s thinking style, providing greater motivation for them to continue answering.
@@ 5.1 Experimental Setups
To fairly evaluate the performance of the PsyDI framework, we have gathered multiple datasets and compared them with both current state-of-the-art open-source and closed-source models. The datasets included in our analysis are Reddit, 16Personalities, and Diamente. Detailed information regarding the datasets, baselines, training settings, and evaluation methods can be found in the Appendix B.
@@ 5.4 Pipeline Evaluation
Thirdly, we also monitor the test interaction procedure of the Character AI bot with PsyDI. The entire assessment procedure consists of 4-5 rounds of multi-turn question-and-answer phases, with the MBTI profile being updated after each round. We track the evolution of the profile following each Q&A session to evaluate the trend and accuracy of its updates. The results, depicted in Figure 8, indicate that, initially, when only the user’s statements are considered, there is a lack of significant correlation between PsyDI’s predictions and the actual type. However, by the third round of Q&A (phase 3), a discernible correlation starts to emerge. By the concluding stage (phase 4), the scores for the true MBTI and the prediction are substantially higher than those for other types, demonstrating PsyDI’s capability to accurately distinguish between various types. Furthermore, due to the bot’s expressions not perfectly reflecting its actual MBTI, PsyDI also assigns higher scores to some similar MBTIs, thereby revealing which types are more closely related; for example, scores for XNFX types, represented by the 4x4 matrix in the top-left corner, tend to cluster together.
@@ 6.1 Psychological measurement
As one of the primary measuring methods in psychology, measurement has been developed a long history and has been established a comprehension measurement system covering different sub-domains, such as clinical diagnosis, military mental state measurement and so on (Cohen et al., 1996). Psychological measurement has many categories, including testing, interview, behavior observation experiment, etc. Testing has been the most preferable tools in many scenery due to its preciseness based on well-designed criterion, consistent measurement despite scenery and convenience in re-usability. However, it needs sufficient expert opinions to design as well as lacks personalized testing to deepen the measurement (Merry et al., 2012). Moreover, expert interview make up these short comes, yet suffers from plenty of resources to collect sparse, non-mass-production-able instances (Crisp & Lincoln, 2014). Although originated from but not belongs to psychological research field, the measurement method of MBTI test has been discussed as having reference value and also gain a wide accept between many individuals (Boyle, 1995; Lamond, 2001).
@@ 6.2 Large Language Models towards Psychological and MBTI
In the advantage of capability emergence firstly evolved out by OpenAI’s GPT models, LLMs could spontaneously capture the deep understanding in contextual relevant contents, achieving human-like level performance (Achiam et al., 2023). Various works have discussed the pros and cons large language models has reached in psychological research field, demonstrate the potential capabilities LLMs mastered in psychological chatting, analysing, guiding and reasoning (Ke et al., 2024; Demszky et al., 2023; Binz & Schulz, 2023). Current cross-over applications includes multi-turn dialogue framework for psychological counseling Zhang et al. (2024), conversational diagnose (Tu et al., 2024), gamified assessments (Yang et al., 2024), introducing more noval pipelines under this scope. Also been a growing force in psycho-cognitive research, LLMs have shown a capable solution to recognize and simulate human cognitive process (Bubeck et al., 2023). This helps MBTI-related AI application exploration whose essence is human cognitive patterns. A series of works have investigated how LLMs captures cognitive personaliies via role-playing different types (Chen et al., 2024; Cui et al., 2023; Tseng et al., 2024; La Cava et al., 2024). LLM-driven measurement abilities are studied to confirm the MBTI awareness also contributes personality discrimination (Rao et al., 2023; Song et al., 2024). A gamified MBTI test has been designed and evaluated on PsychoGAT (Yang et al., 2024).
7 Conclusions and Limitations
@@ Acknowledgments
We thank several members in Shanghai AI Laboratory for their invaluable help and support, which have significantly enhanced the quality of this work. We are thankful to Qingzi Zhu, whose exceptional skills in graphic design were pivotal in creating the visual materials for both the online presentation and this manuscript. Additionally, we acknowledge Mengjie Jiang for her contributions in not only statistically analyzing the experimental data and crafting the key overview figure but also generalizing our framework into another psychology field. Their collective contributions have been indispensable, and we are deeply appreciative of their support.
@@ B.4 Evaluating Method
. To evaluate the score model in PsyDI, for each pair-wise data sample $(p_{i},p_{j},mbti)$, we asked all baseline LLMs to determine which statement was more likely from a user with the specified MBTI type $mbti$. The prompt was designed as follows:

### v4

@@ Abstract
In the field of psychology, traditional assessment methods, such as standardized scales, are frequently critiqued for their static nature, lack of personalization, and reduced participant engagement, while counseling evaluations are often inaccessible to the general public. The complexity of quantifying psychological traits further limits these methods. Despite advances with large language models (LLMs), many still depend on single-round Question-and-Answer interactions. To bridge this gap, we introduce PsyDI, a personalized and progressively in-depth chatbot designed for psychological measurements, exemplified by its application in the Myers-Briggs Type Indicator (MBTI) framework. PsyDI leverages user-related multi-modal information and engages in customized, multi-turn interactions to provide personalized, easily accessible measurements, while ensuring precise MBTI type determination. To address the challenge of unquantifiable psychological traits, we introduce a novel training paradigm that involves learning the ranking of proxy variables associated with these traits, culminating in a robust score model for MBTI measurements. The score model enables PsyDI to conduct comprehensive and precise measurements through multi-turn interactions within a unified estimation context. Through various experiments, we validate the efficacy of both the score model and the PsyDI pipeline, demonstrating its potential to serve as a general framework for psychological measurements. Furthermore, the online deployment of PsyDI has garnered substantial user engagement, with over 5,000 visits, resulting in the collection of numerous multi-turn dialogues annotated with MBTI types, which facilitates further research. The source code for the training and web service components is publicly available as a part of OpenDILab at: https://github.com/opendilab/PsyDI
@@ 2 Problem Formulation
We initiate our analysis by modeling the measurement of psychological indicators in Section 2.1 based on the characteristics of MBTI, as detailed in the Appendix A. Then we introduce the relationship between the MDP modeling and our designed framework in Section 2.2
2.1 MDP Definition for psychological measurement
We conceptualize the interaction between LLM agents and the test taker (user) as a Markov Decision Process (MDP) (Sutton & Barto, 2018). This formulation allows us to systematically optimize the sequence of inquiries posed to the user, with the aim of enhancing the accuracy, efficiency and engagement of MBTI evaluations. Formally, the MDP is articulated as =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\}$. The states $s$ in the observable state space $\mathcal{S}$ are defined as $s=\{p_{1},\ldots,p_{n}\}$, where $p_{i}$ is the $i$-th textual statement of the user (e.g. I like to talk to different people and share my stories). Additionally, the state space includes a special terminated state $s_{\text{exit}}$ representing the user’s decision to exit the evaluation process.
@@ 2.1 MDP Definition for psychological measurement
2.2 Relationship between MDP and Practical Framework
The MDP environment previously defined is an idealized model, providing a clear framework for subsequent analysis. Realistically, we can hardly construct a perfectly ideal MDP due to the inability to directly obtain users’ true psychological indicators, thus precluding accurate reward determination. Instead, we employed a heuristic algorithm to handle reward uncertainty flexibly.
The original action space in MDP we defined is $\mathcal{A}=\mathcal{A}_{q}\cup\mathcal{A}_{t}$, where $\mathcal{A}_{q}$ represents the entirety of possible questions that can be posed. Considering the vastness of the original action space, we propose a heuristic algorithm that leverages an intuitive understanding that a well-formed question should revolve around the user’s statements (Hill et al., 1992). By decomposing the action space into two sequential phases—first selecting a statement and subsequently generating a question centered on that statement—we effectively reduce the dimensionality of the action space. Specifically, the decomposed action space is $\bigcup^{n}_{i=1}(a_{p_{i}}\times Q_{p_{i}})$, where $a_{p_{i}}$ represents for selecting statement $p_{i}$ to ask following question and $Q_{p_{i}}$ is the question space related to statement $p_{i}$.
Leveraging the powerful language capabilities and goal-oriented abilities of LLMs within short conversational rounds, we further extend the question-generating process, denoted as $q_{p_{i}}$, into a multi-turn dialogue. This multi-turn dialogue is a sub-process, where each question is based on the statement $p_{i}$ and the history of question-answer pairs $h_{j}$ within this multi-turn dialogue. Consequently, the action $q_{p_{i}j}$ represents the $j$-th question in the multi-turn dialogue centered around $p_{i}$. Therefore, the action space is $\bigcup^{n}_{i=1}(a_{p_{i}}\times\bigcup^{m}_{j=1}q_{p_{i}j})$. In the subsequent pipeline design, we will demonstrate how to step-by-step select actions within this action space to optimize the return.
@@ 3 PsyDI Progressively In-depth Pipeline
In this section, we introduce the PsyDI pipeline, designed to enhance psychological assessments through a conversational approach. By decomposing the chat process into multi-turn conversations focused on specific statement, PsyDI ensures that large language models (LLMs) stay on track with their instructions. Inspired by traditional psychometric practices, we introduce the MBTI Profile as a continuous context to guide LLM questioning, detailed in Section 3.1. This Profile dynamically updates based on user statements, ensuring the conversation remains psychologically insightful. Additionally, during the multi-turn questioning phase, questions are centered around these informative statements, and a hybrid response format combining multiple-choice and free-form answers enhances user engagement and flexibility, as detailed in Section 3.3. Additionally, a hybrid response format combining multiple-choice and free-form answers enhances user engagement and flexibility, detailed in Section 3.3. Through these integrated strategies, PsyDI aims to provide a robust framework for psychologically informed conversations. The entire workflow of PsyDI is illustrated in Figure 2. For more details on the pipeline, please refer to Appendix D.
@@ 3.3 Multi-turn Chat
Although LLMs perform exceptionally well in single-round, goal-oriented dialogues, it encounters difficulties in maintaining a coherent focus around a single objective when faced with a dozen rounds of multiple questions. To address this limitation, we have devised a hybrid approach that leverages LLMs’ strengths by structuring interactions into modules, each centered on a specific statement. Within these modules, their capabilities are employed to progressively delve into questions related to MBTI, ensuring a step-by-step deepening of the exploration. Specifically, each module begins with a statement from the user, and PsyDI asks 3-5 rounds of progressively deeper questions related to that statement. The user’s responses are then summarized into a single, refined self-description. For example, if the initial statement is "I feel stressed at work," after 3-5 rounds of questioning about the sources of stress, the summarized statement might be: "I feel stressed at work when I cannot meet deadlines. When this happens, I often relieve the stress by reorganizing my plans." The summarization of each module’s dialogue then contributes new insights to the pipeline, facilitating a cumulative and systematic advancement of the overall structure.
@@
For the first challenge, to lead users to reveal their underlying motivations, we utilize multiple-choice questions where each option corresponds to a specific cognitive preference. By presenting such options, users can select the one that best aligns with their own, thereby avoiding ambiguous expressions and implicitly guiding them to reflect on their cognitive tendencies. For questions requiring confirmation, PsyDI employs forced-choice questions, compelling users to identify their most likely cognitive preferences. Also, PsyDI accepts free-form answers, allowing users to provide responses beyond the preset options. The prompt for this phase is list in Appendix C. For the second challenge, which is to keep user’s interesting. We have designed a multimodal input system that allows users to input their preferred music and images, which are then converted into text to form a user statement. Additionally, our questioning framework is designed with a three-phase thought chain. First, we analyze the user statement, then we infer the user’s underlying thought patterns, and finally, we derive the next question based on these patterns. This approach ensures that the entire questioning process revolves around the user’s thinking style, providing greater motivation for them to continue answering. For a clearer example of the full conversation, please refer to Figure 13,  14,  15, and  16.
@@ 4 Score Model
To guide the direction of multi-turn chat using the MBTI Profile, it is essential to first determine the current MBTI estimation of the user. Consequently, we need to train a score model to infer the user’s MBTI from their statements. However, the psychological indicators embedded in these expressions are inherently difficult to observe, and existing datasets of psychological indicators are often noisy and biased. To address these challenges, we propose a method that employs a proxy variable to indirectly observe the original variable, as introduced in Section 4.1. By utilizing rank loss in Section 4.2, the score model focus on learning the rank of the proxy variable rather than its absolute values, thereby ensuring the robustness of score model’s indirect learning of the original variable. We also introduce the pair-wise dataset construction method in Section 4.3. This method generates a dataset where each sample consists of a pair of statements and their corresponding MBTI labels, which is used to train the rank loss.
@@ 5 Experiments
Furthermore, we aim to demonstrate that PsyDI is a generalizable framework that is applicable to a wide range of psychology measurements, not limited to MBTI prediction. We applied PsyDI to two other psychological indicators, Big Five and PANAS-X, to illustrate its effectiveness, as detailed in Appendix E.4. The specific testing process is depicted in Figure 20,  21, and  22.
@@ 5.1 Experimental Setups
To fairly evaluate the performance of the PsyDI framework, we have gathered multiple datasets and compared them with both current state-of-the-art open-source and closed-source models. The datasets included in our analysis are Reddit, 16Personalities, and Diamente. Detailed information regarding the datasets, baselines, training settings, and evaluation methods can be found in the Appendix B, with the distribution of the dataset illustrated in Figure 9.
@@ 5.4 Pipeline Evaluation
Thirdly, we also monitor the test interaction procedure of the Character AI bot with PsyDI. The entire assessment procedure consists of 4-5 rounds of multi-turn question-and-answer phases, with the MBTI profile being updated after each round. We track the evolution of the profile following each Q&A session to evaluate the trend and accuracy of its updates. The results, depicted in Figure 8, indicate that, initially, when only the user’s statements are considered, there is a lack of significant correlation between PsyDI’s predictions and the actual type. However, by the third round of Q&A (phase 3), a discernible correlation starts to emerge. By the concluding phase (phase 4), the scores for the true MBTI and the prediction are substantially higher than those for other types, demonstrating PsyDI’s capability to accurately distinguish between various types. Furthermore, due to the bot’s expressions not perfectly reflecting its actual MBTI, PsyDI also assigns higher scores to some similar MBTIs, thereby revealing which types are more closely related; for example, scores for XNFX types, represented by the 4x4 matrix in the top-left corner, tend to cluster together.
@@ 6.1 Psychological measurement
As one of the primary measuring methods in psychology, measurement has been developed a long history and has been established a comprehension measurement system covering different sub-domains, such as clinical diagnosis, military mental state measurement and so on (Cohen et al., 1996). Psychological measurement has many categories, including testing, interview, behavior observation experiment, etc. Testing has been the most preferable tools in many scenery due to its preciseness based on well-designed criterion, consistent measurement despite scenery and convenience in re-usability. However, it needs sufficient expert opinions to design as well as lacks personalized testing to deepen the measurement (Merry et al., 2012). Moreover, expert interview make up these short comes, yet suffers from plenty of resources to collect sparse, non-mass-production-able instances (Crisp & Lincoln, 2014). The Myers-briggs type indicator (MBTI) is a self-report questionnaire on personality types in personnel management (Myers et al., 1962). This test and current concepts are originated from Carl Jung’s personality types, a typology theory building and investigating an individual’s mental cognitive model (Jung, 1923) and has been extended into 16 personalities by John Beebe (Jung & Beebe, 2016). Although originated from but not belongs to psychological research field, the measurement method of MBTI test has been discussed as having reference value and also gain a wide accept between many individuals (Boyle, 1995; Lamond, 2001). For example, previous research has shown a significant positive correlation between MBTI and the Big Five personality test (Furnham, 1996; 2022).
@@ 6.2 Large Language Models towards Psychological and MBTI
In the advantage of capability emergence firstly evolved out by OpenAI’s GPT models, LLMs could spontaneously capture the deep understanding in contextual relevant contents, achieving human-like level performance (Achiam et al., 2023). Various works have discussed the pros and cons large language models has reached in psychological research field, demonstrate the potential capabilities LLMs mastered in psychological chatting, analysing, guiding and reasoning (Ke et al., 2024; Demszky et al., 2023; Binz & Schulz, 2023). Current cross-over applications includes multi-turn dialogue framework for psychological counseling Zhang et al. (2024), conversational diagnose (Tu et al., 2024), gamified assessments (Yang et al., 2024), introducing more noval pipelines under this scope. Also been a growing force in psycho-cognitive research, LLMs have shown a capable solution to recognize and simulate human cognitive process (Bubeck et al., 2023). This helps MBTI-related AI application exploration whose essence is human cognitive patterns. A series of works have investigated how LLMs captures cognitive personaliies via role-playing different types (Chen et al., 2024; Cui et al., 2023; Tseng et al., 2024; La Cava et al., 2024; Kwan et al., 2024). LLM-driven measurement abilities are studied to confirm the MBTI awareness also contributes personality discrimination (Rao et al., 2023; Song et al., 2024). A gamified MBTI test has been designed and evaluated on PsychoGAT (Yang et al., 2024).
7 Discussion
We have included safety measures in the prompts of LLMs to ensure they remain secure. Specifically, we ask the LLM to check if a user’s input is aggressive or manipulative. If it is, the LLM stops and asks for a new, safe question. This helps keep the conversation safe. However, we need to continuously improve this system to catch all potential issues.
We have added a privacy page in PsyDI Online to inform users about how their data will be used. We also ensure all data is anonymized. This helps protect user privacy. However, we must follow new privacy regulations and continue to improve our methods to maintain high standards.
It is important to acknowledge that while PsyDI has shown promising results, it has not yet undergone rigorous reliability testing. Before this framework can be fully adopted in real-world applications, it is imperative to conduct extensive reliability testing.
8 Conclusions
@@ Acknowledgments
We thank several members in Shanghai AI Laboratory for their invaluable help and support, which have significantly enhanced the quality of this work. We are thankful to Qingzi Zhu, whose exceptional skills in graphic design were pivotal in creating the visual materials for both the online presentation and this manuscript. Additionally, we acknowledge Mengjie Jiang for her contributions in not only statistically analyzing the experimental data and crafting the key overview figure but also generalizing our framework into another psychology field. Their collective contributions have been indispensable, and we are deeply appreciative of their support. Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper. Only add this information once your submission is accepted and deanonymized.
@@ B.1 Datasets
During the dataset construction process, we carefully managed the distribution and quality of the data. Specifically, we filtered out statements based on their length and sampled 400 instances for each MBTI type, ensuring a balanced and representative dataset.
We also examined the length distribution and topic coverage of the dataset. For example, using the Reddit dataset, we created a box plot in Figure 9(a) to show the length distribution for all MBTI types. The plot shows that the length distributions are quite similar across all MBTI types, with most texts ranging from 100 to 300 words, although some longer texts are also present. This indicates that the dataset exhibits a consistent text length across different MBTI types, allowing the model to focus more on the personality information contained within balanced text lengths In Figure 9(b), we displayed the main verbs and their corresponding verb-noun phrases in the dataset. This analysis shows that the dataset covers a wide range of topics and includes many verb-noun phrases strongly related to MBTI, such as "make decision."
(a) Box plot of length distribution of Reddit dataset
(b) Top verb-noun phrases in Reddit dataset
Figure 9: Distribution of Reddit dataset
@@ B.4 Evaluating Method
To evaluate the score model in PsyDI, for each pair-wise data sample $(p_{i},p_{j},mbti)$, we asked all baseline LLMs to determine which statement was more likely from a user with the specified MBTI type $mbti$. The prompt was designed as follows:
@@
Furthermore, we assess the accuracy of MBTI bot in reflecting the most typical options compared to human responses, to determine whether MBTI bots like Characeter AI can be effectively used as testers. We use INTP and ESFP as examples and select several topics to generate related questions with options aligned with NF/NT/SP/SJ dimensions. We compare the accuracy of bot and human responses to these questions.
As illustrated in Figure 10, under most topics, the bot achieves a level of accuracy that is comparable to, or even higher than, that of humans. This indicates that the bot can simulate responses typical of a specific personality type, and even more representative of typical traits compared to individual humans. This finding supports the validity of using bots as substitutes for human participants as tester.
(a) Bot vs Human on MBTI Related Question (INTP)
(b) Bot vs Human on MBTI Related Question (ESFP)
Figure 10: Accuracy of MBTI Bot vs. Human Responses Across Various Topics. For each MBTI-related topic, PsyDI generates a question stem and four options aligned with NF/NT/SP/SJ dimensions. Both the MBTI bot and human participants are asked to select one option. Under most topics, the bot achieves a level of accuracy that is comparable to, or even higher than, that of humans.